OpenStudy (anonymous):

15. A ball is thrown from ground level with a velocity of 24.5 meters per second. As it starts back down to the ground, it hits a tree at a point 29.4 meters from the ground. How many seconds is the ball in the air before it hits the tree? Use h = vt – 4.9t2 in your work. A. 1 B. 2 C. 3 D. 5

OpenStudy (anonymous):

I solved the problem and I got 4.9(t^2 - 5t + 6) and for t I got 2 seconds or 3 seconds But I can't have both as the answer

OpenStudy (anonymous):

Can you please help me

OpenStudy (anonymous):

2 seconds

OpenStudy (anonymous):

why

OpenStudy (anonymous):

plug it into the equation they gave you

OpenStudy (anonymous):

But I also plugged in 3 and It works

OpenStudy (anonymous):

29.4= 24.5(2)– 4.9(2)^2 29.4=49-19.6 29.4=29.4

OpenStudy (anonymous):

oh wait 3 does work huh...

OpenStudy (anonymous):

29.4 = 24.5(3) -4.9(3)^2 29.4 = 73.5 - 44.1 29.4 = 29.4

OpenStudy (anonymous):

That's why I am confused

OpenStudy (anonymous):

did you ask your teacher?

OpenStudy (anonymous):

Well I am doing this online Mizzou University Subject so I don't have anyone to ask But still I think the answer is 3 seconds because it says that it goes at 24.5 meters per second so it would just take it 2 seconds to go up and down and it meets the 29.4 meters s o I think it will be 3 seconds

OpenStudy (anonymous):

Yeah that's what I was thinking, this isnt physics right?

OpenStudy (anonymous):

I will put 3 as the answer and I will tell you if I got it right in 2 seconds ok

OpenStudy (anonymous):

what do you say, should I do it or not

OpenStudy (anonymous):

yep do it

OpenStudy (anonymous):

Everything got erased I don't know how so I will do this tonight and If you still want to know if I got it I will post it on this

OpenStudy (anonymous):

A;right :) I do want to know lol

OpenStudy (anonymous):

Hey I did everything all over again and I just submitted it and I got that problem right woo hoo! :D

OpenStudy (anonymous):

Thanks man. Thanks for everthing.

OpenStudy (anonymous):

*everything

OpenStudy (anonymous):

I guess I will see you later. Bye